OpenStudy (anonymous):

How would I show my work to make this true? x = -5 , x = 2 x^2 +3x -10 =0

OpenStudy (anonymous):

its asking if x=-5 and x=2 are solutions to the equation x^2+3x-10=0 to prove this is true, factor the given equation (by whatever method you want) and you will arrive at the factored equation (x+5)*(x-2)=0, which shoes solutions of x=-5, 2, proving given data

OpenStudy (anonymous):

shows*

OpenStudy (anonymous):

So is this enough steps shown, or am I missing steps to prove the answer? (x+5)*(x-2)=0 x^2 +3x -10 =0

OpenStudy (anonymous):

my bad..

OpenStudy (anonymous):

for what?

OpenStudy (anonymous):

that would be plenty for any teacher i have ever had

OpenStudy (anonymous):

Oh, I just thought they were expressions not equations... I to concentrate a little more..

OpenStudy (anonymous):

Sweet. I just wanted to make sure I had it right.

OpenStudy (anonymous):

if you really want to show more work, you can use the quadratic equation to validate your a and b values in the factored expression, but idk if your prof would really require that

OpenStudy (anonymous):

Idk either. My teacher is strict and constantly picks at me lol

OpenStudy (anonymous):

if thats the case, i would recommend the quadratic equation, its easy and keeps profs off your back when they want to be mean

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

That's what it's suppose to be is a quadratic equation. I wasn't thinking

OpenStudy (anonymous):

for an equation of the form Ax^2+Bx+C=0, the quadratic equation is \[(-B \pm \sqrt{B^2-4*A*C})/(2*A)\]

OpenStudy (anonymous):

so B=3, A=1, C=-10, solving it should get you 2 and -5 really easily

OpenStudy (anonymous):

It looks like a foreign language lol Unfortunately, I have to try and re-learn all this every single day.

OpenStudy (anonymous):

ill walk you through it if you want

OpenStudy (anonymous):

That would be awesome!

OpenStudy (anonymous):

A=1, B=3, C=-10 so (−B±√B2−4∗A∗C)/(2∗A) becomes (-3±√ (3^2-4*1*-10))/(2*1) (-3±√(9+40))/2 (-3±√49)/2 (-3±7)/2 this gets you two answers out of the + and the -, for the -, it becomes (-3-7)/2 -10/2 -5 for the + you get (-3+7)/2 4/2 2 so your answers are -5 and 2

OpenStudy (anonymous):

well, x=-5 and x=2 if you want to be explicit

OpenStudy (anonymous):

Wow! Thank you so much! I wish I could give you another medal :)

OpenStudy (anonymous):

thanks, i really enjoy helping out ^^

OpenStudy (anonymous):

:)